1 Find the Taylor series for notation. f(x) at C = 4. Write the Taylor series in sigma х
Answer #1

oluer We know that the Taylors series of a function fro is given by •W fa flas (x-aja Ko K! Here f(n) f(X) a = 4 2 Now 7 °(+1) = - 120 +14) = (-1515!) 45+1 x f(x1 720 = f(4) = x? 16 (6!) 40+1 #*(X) = Guk k! => +*(4) - Cukki akti (4)* K + Thus **(4Please feel free to ask any query in the comment box and don't forget to rate if you like